A PHY-102 student at Mulungushi University wants to make a simple 10 𝑛𝐹 capacitor for tuning her crystal radio, using two sheets of aluminum foil as plates, with a few sheets of paper between them as a dielectric. The paper has a dielectric constant of 3.0, and the thickness of one sheet of it is 0.20 mm. If the sheets of paper measure 22Γ—28 π‘π‘š and she cuts the aluminum foil to the same dimensions, how many sheets of paper would she use between her plates to get the proper capacitance?


1
Expert's answer
2021-05-03T10:31:42-0400

Capacitance, "C=10\space nF=10\times10^{-9}\space F"

Dielectric constant, "K=3"

Thickness of one sheet, "d=0.20\space mm=0.20\times10^{-3}\space m"

Area of plates, "A=22\times28\space cm^2=616\times10^{-4}\space m^2"

"C_o=\dfrac{K\epsilon_o A}{d}"

"C_o=\dfrac{3\times8.85\times10^{-12}\times616\times10^{-4}}{0.20\times10^{-3}}"

"C_o=8.1774\times10^{-9}\space F"

Number of sheets of paper used be "n"

Capacitance needed, "C=10\times10^{-9}\space F"

Equivalent capacitance,

"\dfrac{1}{C}=\bigg[\dfrac{1}{C_o}+\dfrac{1}{C_o}+\dfrac{1}{C_o}+........+\dfrac{1}{C_o}\bigg]"

"C=\dfrac{C_o}{n}"

"n=\dfrac{C_o}{C}=\dfrac{8.1774\times10^{-9}}{10\times10^{-9}}\approx1"